OpenStudy (anonymous):

i need help with this 8^2x=11 solve the giver equation

OpenStudy (solomonzelman):

what comes to mind, is taking log(base 8) of both sides.

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{is it } \hspace{.33em}\\~\\ 8^2x=11 \hspace{.33em}\\~\\ \normalsize \text{or} \hspace{.33em}\\~\\ 8^{2x}=11 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

8^2x=11 so the secant one

OpenStudy (mathmath333):

u mean this one \(\LARGE \begin{align} \color{black}{ 8^{2x}=11 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

OK

OpenStudy (anonymous):

ya i don't under stand how to do it

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 8^{2x}=11 \hspace{.33em}\\~\\ (8^{2})^x=11 \hspace{.33em}\\~\\ (64)^x=11 \hspace{.33em}\\~\\ \log (64)^x=\log 11 \hspace{.33em}\\~\\ x\log (64)=\log 11 \hspace{.33em}\\~\\ x=\dfrac{\log 11}{\log 64} \hspace{.33em}\\~\\ }\end{align}\) see it makes sense

OpenStudy (mathmath333):

see if it makes sense

OpenStudy (anonymous):

ok but it is not telling me what the ancer

OpenStudy (anonymous):

2.0794 c. 2.3979 0.5766 d. 6.0902 it is one of them

OpenStudy (mathmath333):

that can be solved further by taking logs . so we have to take log for that.

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

have they given log table ?

OpenStudy (anonymous):

no

OpenStudy (mathmath333):

we can use the wolfram http://www.wolframalpha.com/input/?i=%5Cdfrac%7B%5Clog+11%7D%7B%5Clog+64%7D%3D

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

so its \(\large \begin{align} \color{black}{ 0.5766 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\[\Huge \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#E6E600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#E6E600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{o}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{s}}\color{#E6E600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{ }}\color{red}{\text{}}\end{array} \]

OpenStudy (anonymous):

ty so much